Let A = {2,4} and B = {6,8,10} and define relations R and S from A to B as follows: For all (x, y) E A ® B, XRy iff x|y, and xSy iff y - 4 = x. State explicitly which ordered pairs are in A x B, R, S, RUS, and RNS


1
Expert's answer
2021-04-15T07:32:36-0400


Let "A = \{2,4\}" and "B = \{6,8,10\}" and define relations "R" and "S" from "A" to "B" as follows:


For all "(x, y) \in A \times B""xRy" iff "x|y", and "xSy" iff "y - 4 = x".


Let us state explicitly which ordered pairs are in "A \times B, R, S, R\cup S," and "R\cap S":


"A\times B=\{(2,6),(2,8),(2,10),(4,6),(4,8),(4,10)\}"


"R=\{(2,6),(2,8),(2,10),(4,8)\}"


"S=\{(2,6),(4,8)\}"


"R\cup S=\{(2,6),(2,8),(2,10),(4,8)\}"


"R\cap S=\{(2,6),(4,8)\}"